OpenStudy (qqstory):

why does -8ln|sqrt(2) -1| = 8(ln(1/sqrt(2)-1) ?

OpenStudy (qqstory):

Can someone explain or prove this? Thanks

OpenStudy (qqstory):

\[-8 \ln \left| \sqrt{2} -1 \right| = 8 \ln \left| \frac{ 1 }{ \sqrt{2} -1 } \right|\]

jimthompson5910 (jim_thompson5910):

\[\Large -8 \ln \left| \sqrt{2} -1 \right| = 8\ln \left| (\sqrt{2} -1)^{-1} \right|\] \[\Large -8 \ln \left| \sqrt{2} -1 \right| = 8 \ln \left| \frac{ 1 }{ \sqrt{2} -1 } \right|\] I used this rule y*ln(x) = ln(x^y) for step 1 Then I used this rule for step 2: x^(-k) = 1/(x^k)